A 1.14 MΩ voltmeter is placed in parallel with a 79.0 kΩ resistor in a circuit.

(b)

What is the resistance of the combination? (Enter your answer in ohms to at least 3 significant figures.)

(c)

If the voltage across the combination is kept the same as it was across the 79.0 kΩ resistor alone, what is the percent increase in current?

(d)

If the current through the combination is kept the same as it was through the 79.0 kΩ resistor alone, what is the percentage decrease in voltage?

Are the changes found in parts (c) and (d) significant? Discuss.

Update: Got everything right except D
Answer #1

D)

Equivalent resistance, Req = (79000 x 1140000)/(79000 + 1140000) = 73.88 k ohm

(V - Vo)/Vo = [(I(73880.23) - I(79000))/I x 79000] x 100 = 6.48 %

Comment in case any doubt please rate my answer...